# VolleyBots: A Testbed for Multi-Drone Volleyball Game Combining Motion Control and Strategic Play

### v3

@@ D.1 Back and Forth
The drone is initialized at an anchor position $(4.5,0,2)$, i.e., the center of the red court with a height of $2\,m$. The other anchor position is $(9.0,4.5,2)$, with the target points switching between two designated anchor positions. The drone is required to sprint between two designated anchors to complete as many round trips as possible. $5$ steps within a sphere with a $0.6\,m$ radius near the anchor position are required for each stay. The maximum episode length is $800$ steps.
@@
This task is evaluated by the number of target points reached within the time limit. A successful stay is defined as the drone staying $5$ steps within a sphere with a $0.6\,m$ radius near the target anchor.
@@ D.2 Hit the Ball
The drone is initialized randomly around an anchor position $(4.5,0,2)$, i.e., the center of the red court with a height of $2\,m$. The drone’s initial position is sampled uniformly random from $[4,-0.5,1.8]$ to $[5,0.5,2.2]$. The ball is initialized at $(4.5,0,5)$, i.e., $3\,m$ above the anchor position. The ball starts with zero velocity and falls freely. The drone is required to perform a single hit to strike the ball toward the opponent’s court, i.e., in the negative direction of the x-axis, aiming for maximum distance. The maximum episode length is $800$ steps.
@@ D.3 Solo Bump
The drone is initialized randomly around an anchor position $(4.5,0,2)$, i.e., the center of the red court with a height of $2\,m$. The drone’s initial position is sampled uniformly random from $[4,-0.5,1.8]$ to $[5,0.5,2.2]$. The ball is initialized at $(4.5,0,4)$, i.e., $2\,m$ above the anchor position. The ball starts with zero velocity and falls freely. The drone is required to stay within a sphere with $1\,m$ radius near the anchor position and bump the ball as many times as possible. A minimum height of $3.5\,m$ is required for each bump. The maximum episode length is $800$ steps.
@@
This task is evaluated by the number of successful consecutive bumps performed by the drone. A successful bump is defined as the drone hitting the ball such that the ball’s highest height exceeds $3.5\,m$ but not exceeds $4.5\,m$.
@@ D.4 Bump and Pass
Drone 1 is initialized randomly around anchor 1 with position $(4.5,-2.5,2)$, and Drone 2 is initialized randomly around anchor 2 with position $(4.5,2.5,2)$. The initial position of drone 1 is sampled uniformly random from $(4,-3,1.8)$ to $(5,-2,2.2)$, and the initial position of drone 2 is sampled uniformly random from $(4,2,1.8)$ to $(5,3,2.2)$. The ball is initialized at $(4.5,-2.5,4)$, i.e., $2\,m$ above anchor 1. The ball starts with zero velocity and falls freely. The drones are required to stay within a sphere with $0.5\,m$ radius near their anchors and bump the ball to pass it to each other in turns as many times as possible. A minimum height of $4\,m$ is required for each bump. The maximum episode length is 800 steps.
@@
This task is evaluated by the number of successful consecutive bumps performed by the drones. A successful bump is defined as the drone hitting the ball such that the ball’s highest height exceeds $4\,m$ and lands near the other drone.
@@ D.5 Set and Spike (Easy)
Drone 1 (setter) is initialized randomly around anchor 1 with position $(2,-2.5,2.5)$, and Drone 2 (attacker) is initialized randomly around anchor 2 with position $(2,2.5,3.5)$. The initial position of drone 1 is sampled uniformly random from $(1.5,-3,2.3)$ to $(2.5,-2,2.7)$, and the initial position of drone 2 is sampled uniformly random from $(1.5,2,3.3)$ to $(2.5,3,3.7)$. The ball is initialized at $(2,-2.5,4.5)$, i.e., $2\,m$ above anchor 1. The ball starts with zero velocity and falls freely. The drones are required to stay within a sphere with $0.5\,m$ radius near their anchors. The setter is required to pass the ball to the attacker, and the attacker then spikes the ball downward to the target region in the opposing side. The target region is a circular area on the ground, centered at $(4.5,0)$ with a radius of $1\,m$. The maximum episode length is 800 steps.
@@
[…]
@@ D.6 Set and Spike (Hard)
Drone 1 (setter) is initialized randomly around anchor 1 with position $(2,-2.5,2.5)$, and Drone 2 (attacker) is initialized randomly around anchor 2 with position $(2,2.5,3.5)$. The initial position of drone 1 is sampled uniformly random from $(1.5,-3,2.3)$ to $(2.5,-2,2.7)$, and the initial position of drone 2 is sampled uniformly random from $(1.5,2,3.3)$ to $(2.5,3,3.7)$. The ball is initialized at $(2,-2.5,4.5)$, i.e., $2\,m$ above anchor 1. The ball starts with zero velocity and falls freely. The racket is initialized at $(-4,0,0.5)$, i.e., the center of the opposing side. The drones are required to stay within a sphere with $0.5\,m$ radius near their anchors. The setter is required to pass the ball to the attacker, and the attacker then spikes the ball downward to the opponent’s court without being intercepted by the defense racket. The maximum episode length is 800 steps.
@@
[…]
@@
Two drones are required to play 1 vs 1 volleyball in a reduced-size court of $6\,m\times 3\,m$. Drone 1 is initialized randomly around anchor 1 with position $(1.5,0.0,2.0)$, i.e., the center of the red court with height $2\,m$, and Drone 2 is initialized randomly around anchor 2 with position $(-1.5,0.0,2.0)$, i.e., the center of the blue court with height $2\,m$. The initial position of drone 1 is sampled uniformly random from $(1.4,-0.1,1.9)$ to $(1.6,0.1,2.1)$, and the initial position of drone 2 is sampled uniformly random from $(-1.4,-0.1,1.9)$ to $(-1.6,0.1,2.1)$. At the start of a game (i.e. an episode), one of the two drones is randomly chosen to serve the ball, which is initialized $1.5\,m$ above the drone. The ball starts with zero velocity and falls freely. The game ends when one of the drones wins the game or one of the drones crashes. The maximum episode length is 800 steps.
@@
The task involves two teams of drones competing in a 3 vs 3 volleyball match within a reduced-size court of $9\,m\times 4.5\,m$. Drone 1, Drone 2, and Drone 3 belong to Team 1 and are initialized at positions $(3.0,-1.5,2.0)$, $(3.0,1.5,2.0)$ and $(6.0,0.0,2.0)$ respectively. Similarly, Drone 4, Drone 5, and Drone 6 belong to Team 2 and are initialized at positions $(-3.0,-1.5,2.0)$, $(-3.0,1.5,2.0)$ and $(-6.0,0.0,2.0)$ respectively. At the start of a game (i.e., an episode), one of the two teams is randomly selected to serve the ball. The ball is initialized at a position $3\,m$ directly above the serving drone. The ball starts with zero velocity and falls freely. The game ends when one of the teams wins the game or one of the drones crashes. The maximum episode length is 500 steps.
@@
The task involves two teams of drones competing in a 6 vs 6 volleyball match within a standard-size court of $12\,m\times 6\,m$. Drone 1 to Drone 6 belong to Team 1 and are initialized at positions $(3.0,-3.0,2.0)$, $(3.0,0.0,2.0)$, $(3.0,3.0,2.0)$, $(6.0,-3.0,2.0)$, $(9.0,0.0,2.0)$ and $(6.0,3.0,2.0)$ respectively. Similarly, Drone 7 to Drone 12 belong to Team 2 and are initialized at positions $(-3.0,3.0,2.0)$, $(-3.0,0.0,2.0)$, $(-3.0,-3.0,2.0)$, $(-6.0,3.0,2.0)$, $(-9.0,0.0,2.0)$ and $(-6.0,-3.0,2.0)$ respectively. At the start of a game (i.e., an episode), one of the two teams is randomly selected to serve the ball. The ball is initialized at a position $3\,m$ directly above the serving drone. The ball starts with zero velocity and falls freely. The game ends when one of the teams wins the game or one of the drones crashes. The maximum episode length is 500 steps.
@@ Appendix F Details of benchmark experiments
gamma $0.95$ tau $0.005$ target update interval $4$
@@
buffer length $64$ num minibatches $16$ ppo epochs $4$
@@
discrete bin $2$ tau $0.005$ target update interval $4$
@@ F.1 Experimental Platform and Computational Resources
Our experiments were conducted on a workstation equipped with NVIDIA GeForce RTX 4090 or RTX 3090 GPUs, 128 GB of RAM, and Ubuntu 20.04 LTS. The software environment included CUDA 12.4, Python 3.10, PyTorch 2.0, and NVIDIA Isaac Sim 2023.1.0. All single-agent and multi-agent cooperative experiments completed in under 12 hours, while multi-agent competitive experiments finished in under 24 hours.
@@ F.3 Results of single-agent tasks
Fig. 6 plots the training progress of five single-agent algorithms on three single-agent volleyball tasks under both CTBR (top row) and PRT (bottom row) action space, averaged over three seeds. Across every task, PPO (orange) converges fastest and to the highest performance, stabilizing at roughly $9-10$ successful reaches in Back and Forth, $10-11$m in Hit the Ball, and $8-12$ bumps in Solo Bump. TD3 (green), SAC (purple), and DDPG (blue) exhibit comparable moderate performance across most tasks, with TD3 notably outperforming SAC and DDPG in Solo Bump. In contrast, DQN (red) fails to make meaningful progress in any of the tasks. Moreover, each algorithm exhibits comparable behavior under both CTBR and PRT action spaces, with slightly better final performance under PRT for most methods and tasks.
Figure 6: Training curves of single-agent tasks over three seeds.
Figure 7: Training curves of multi-agent cooperative tasks over three seeds.

### v6

@@ D.1 Back and Forth
The drone is initialized at an anchor position $(4.5,0,2)$, i.e., the center of the red court with a height of 2m2\,m. The other anchor position is $(9.0,4.5,2)$, with the target points switching between two designated anchor positions. The drone is required to sprint between two designated anchors to complete as many round trips as possible. $5$ steps within a sphere with a 0.6m0.6\,m radius near the anchor position are required for each stay. The maximum episode length is $800$ steps.
@@
This task is evaluated by the number of target points reached within the time limit. A successful stay is defined as the drone staying $5$ steps within a sphere with a 0.6m0.6\,m radius near the target anchor.
@@ D.2 Hit the Ball
The drone is initialized randomly around an anchor position $(4.5,0,2)$, i.e., the center of the red court with a height of 2m2\,m. The drone’s initial position is sampled uniformly random from [4,0.5,1.8][4,-0.5,1.8] to $[5,0.5,2.2]$. The ball is initialized at $(4.5,0,5)$, i.e., 3m3\,m above the anchor position. The ball starts with zero velocity and falls freely. The drone is required to perform a single hit to strike the ball toward the opponent’s court, i.e., in the negative direction of the x-axis, aiming for maximum distance. The maximum episode length is $800$ steps.
@@ D.3 Solo Bump
The drone is initialized randomly around an anchor position $(4.5,0,2)$, i.e., the center of the red court with a height of 2m2\,m. The drone’s initial position is sampled uniformly random from [4,0.5,1.8][4,-0.5,1.8] to $[5,0.5,2.2]$. The ball is initialized at $(4.5,0,4)$, i.e., 2m2\,m above the anchor position. The ball starts with zero velocity and falls freely. The drone is required to stay within a sphere with 1m1\,m radius near the anchor position and bump the ball as many times as possible. A minimum height of 3.5m3.5\,m is required for each bump. The maximum episode length is $800$ steps.
@@
This task is evaluated by the number of successful consecutive bumps performed by the drone. A successful bump is defined as the drone hitting the ball such that the ball’s highest height exceeds 3.5m3.5\,m but not exceeds 4.5m4.5\,m.
@@ D.4 Bump and Pass
Drone 1 is initialized randomly around anchor 1 with position (4.5,2.5,2)(4.5,-2.5,2), and Drone 2 is initialized randomly around anchor 2 with position $(4.5,2.5,2)$. The initial position of drone 1 is sampled uniformly random from (4,3,1.8)(4,-3,1.8) to (5,2,2.2)(5,-2,2.2), and the initial position of drone 2 is sampled uniformly random from $(4,2,1.8)$ to $(5,3,2.2)$. The ball is initialized at (4.5,2.5,4)(4.5,-2.5,4), i.e., 2m2\,m above anchor 1. The ball starts with zero velocity and falls freely. The drones are required to stay within a sphere with 0.5m0.5\,m radius near their anchors and bump the ball to pass it to each other in turns as many times as possible. A minimum height of 4m4\,m is required for each bump. The maximum episode length is 800 steps.
@@
This task is evaluated by the number of successful consecutive bumps performed by the drones. A successful bump is defined as the drone hitting the ball such that the ball’s highest height exceeds 4m4\,m and lands near the other drone.
@@ D.5 Set and Spike (Easy)
Drone 1 (setter) is initialized randomly around anchor 1 with position (2,2.5,2.5)(2,-2.5,2.5), and Drone 2 (attacker) is initialized randomly around anchor 2 with position $(2,2.5,3.5)$. The initial position of drone 1 is sampled uniformly random from (1.5,3,2.3)(1.5,-3,2.3) to (2.5,2,2.7)(2.5,-2,2.7), and the initial position of drone 2 is sampled uniformly random from $(1.5,2,3.3)$ to $(2.5,3,3.7)$. The ball is initialized at (2,2.5,4.5)(2,-2.5,4.5), i.e., 2m2\,m above anchor 1. The ball starts with zero velocity and falls freely. The drones are required to stay within a sphere with 0.5m0.5\,m radius near their anchors. The setter is required to pass the ball to the attacker, and the attacker then spikes the ball downward to the target region in the opposing side. The target region is a circular area on the ground, centered at $(4.5,0)$ with a radius of 1m1\,m. The maximum episode length is 800 steps.
@@
This task is evaluated by the success rate of set and spike. A successful set and spike consist of four parts, (1) setter_hit: the setter hits the ball; (2) attacker_hit: the attacker hits the ball; (3) downward_spike: the velocity of the ball after the attacker hit is downward, i.e., $v_{z}<0$; (4) in_target: the ball’s landing position is within the target region. The success rate is computed as $1/4\times(\text{setter\_hit}+\text{attacker\_hit}+\text{downward\_spike}+\text{in\_target})$.
@@ D.6 Set and Spike (Hard)
Drone 1 (setter) is initialized randomly around anchor 1 with position (2,2.5,2.5)(2,-2.5,2.5), and Drone 2 (attacker) is initialized randomly around anchor 2 with position $(2,2.5,3.5)$. The initial position of drone 1 is sampled uniformly random from (1.5,3,2.3)(1.5,-3,2.3) to (2.5,2,2.7)(2.5,-2,2.7), and the initial position of drone 2 is sampled uniformly random from $(1.5,2,3.3)$ to $(2.5,3,3.7)$. The ball is initialized at (2,2.5,4.5)(2,-2.5,4.5), i.e., 2m2\,m above anchor 1. The ball starts with zero velocity and falls freely. The racket is initialized at (4,0,0.5)(-4,0,0.5), i.e., the center of the opposing side. The drones are required to stay within a sphere with 0.5m0.5\,m radius near their anchors. The setter is required to pass the ball to the attacker, and the attacker then spikes the ball downward to the opponent’s court without being intercepted by the defense racket. The maximum episode length is 800 steps.
@@
This task is evaluated by the success rate of set and spike. A successful set and spike consist of four parts, (1) setter_hit: the setter hits the ball; (2) attacker_hit: the attacker hits the ball; (3) downward_spike: the velocity of the ball after the attacker hit is downward, i.e., $v_{z}<0$; (4) success_spike: the ball’s landing position is within the opponent’s court without being intercepted by the defense racket. The success rate is computed as $1/4\times(\text{setter\_hit}+\text{attacker\_hit}+\text{downward\_spike}+\text{success\_spike})$.
@@
Two drones are required to play 1 vs 1 volleyball in a reduced-size court of $6\,m\times 3\,m$. Drone 1 is initialized randomly around anchor 1 with position $(1.5,0.0,2.0)$, i.e., the center of the red court with height 2m2\,m, and Drone 2 is initialized randomly around anchor 2 with position (1.5,0.0,2.0)(-1.5,0.0,2.0), i.e., the center of the blue court with height 2m2\,m. The initial position of drone 1 is sampled uniformly random from (1.4,0.1,1.9)(1.4,-0.1,1.9) to $(1.6,0.1,2.1)$, and the initial position of drone 2 is sampled uniformly random from (1.4,0.1,1.9)(-1.4,-0.1,1.9) to (1.6,0.1,2.1)(-1.6,0.1,2.1). At the start of a game (i.e. an episode), one of the two drones is randomly chosen to serve the ball, which is initialized 1.5m1.5\,m above the drone. The ball starts with zero velocity and falls freely. The game ends when one of the drones wins the game or one of the drones crashes. The maximum episode length is 800 steps.
@@
The task involves two teams of drones competing in a 3 vs 3 volleyball match within a reduced-size court of $9\,m\times 4.5\,m$. Drone 1, Drone 2, and Drone 3 belong to Team 1 and are initialized at positions (3.0,1.5,2.0)(3.0,-1.5,2.0), $(3.0,1.5,2.0)$ and $(6.0,0.0,2.0)$ respectively. Similarly, Drone 4, Drone 5, and Drone 6 belong to Team 2 and are initialized at positions (3.0,1.5,2.0)(-3.0,-1.5,2.0), (3.0,1.5,2.0)(-3.0,1.5,2.0) and (6.0,0.0,2.0)(-6.0,0.0,2.0) respectively. At the start of a game (i.e., an episode), one of the two teams is randomly selected to serve the ball. The ball is initialized at a position 3m3\,m directly above the serving drone. The ball starts with zero velocity and falls freely. The game ends when one of the teams wins the game or one of the drones crashes. The maximum episode length is 500 steps.
@@
The task involves two teams of drones competing in a 6 vs 6 volleyball match within a standard-size court of $12\,m\times 6\,m$. Drone 1 to Drone 6 belong to Team 1 and are initialized at positions (3.0,3.0,2.0)(3.0,-3.0,2.0), $(3.0,0.0,2.0)$, $(3.0,3.0,2.0)$, (6.0,3.0,2.0)(6.0,-3.0,2.0), $(9.0,0.0,2.0)$ and $(6.0,3.0,2.0)$ respectively. Similarly, Drone 7 to Drone 12 belong to Team 2 and are initialized at positions (3.0,3.0,2.0)(-3.0,3.0,2.0), (3.0,0.0,2.0)(-3.0,0.0,2.0), (3.0,3.0,2.0)(-3.0,-3.0,2.0), (6.0,3.0,2.0)(-6.0,3.0,2.0), (9.0,0.0,2.0)(-9.0,0.0,2.0) and (6.0,3.0,2.0)(-6.0,-3.0,2.0) respectively. At the start of a game (i.e., an episode), one of the two teams is randomly selected to serve the ball. The ball is initialized at a position 3m3\,m directly above the serving drone. The ball starts with zero velocity and falls freely. The game ends when one of the teams wins the game or one of the drones crashes. The maximum episode length is 500 steps.
@@ Appendix F Details of benchmark experiments
gamma $0.95$ tau $0.005$ target update interval 44
@@
buffer length $64$ num minibatches $16$ ppo epochs 44
@@
discrete bin 22 tau $0.005$ target update interval 44
@@ F.1 Experimental Platform and Computational Resources
Our experiments were conducted on a workstation equipped with NVIDIA GeForce RTX 4090 or RTX 3090 GPUs, 128 GB of RAM, and Ubuntu 20.04 LTS. The software environment included CUDA 12.4, Python 3.10, PyTorch 2.0, and NVIDIA Isaac Sim 2023.1.0. All single‐agent and multi‐agent cooperative experiments completed in under 12 hours, while multi‐agent competitive experiments finished in under 24 hours.
@@ F.2 Hyperparameters of benchmarking algorithms
In the single-agent setting, we tune hyperparameters on a simpler task, Hover, proposed in OmniDrones [41]. In this task, the drone starts from a randomized position and heading, moves toward a randomized target, and then maintains a stable pose without drift. The reward function depends on position error, heading alignment, uprightness, and angular stability. We perform random search over the hyperparameter space, and the best configurations found on Hover yield the baseline performance shown in Table 19. On this simple task, DDPG, TD3, SAC, and PPO achieve comparable performance, while DQN fails. For fairness and reproducibility, we fix each algorithm’s configuration obtained on Hover and apply it unchanged to all other single-agent tasks reported in Table 2. This setup allows us to evaluate cross-task robustness without task-specific tuning.
Table 19: Results of hyperparameter tuning for single-agent RL algorithms on the Hover task.
Algorithms DDPG TD3 SAC PPO
Return $1168.01\pm 16.83$ $1212.82\pm 8.81$ $249.39\pm 4.55 11$$196.68\pm 3.01$
@@ F.3 Results of single-agent tasks
Fig. 7 plots the training progress of five single‐agent algorithms on three single‐agent volleyball tasks under both CTBR (top row) and PRT (bottom row) action space, averaged over five seeds. Across every task, PPO (orange) converges fastest and to the highest performance, stabilizing at roughly 9109-10 successful reaches in Back and Forth, 101110-11m in Hit the Ball, and 8128-12 bumps in Solo Bump. TD3 (green), SAC (purple), and DDPG (blue) exhibit comparable moderate performance across most tasks, with TD3 notably outperforming SAC and DDPG in Solo Bump. In contrast, DQN (red) fails to make meaningful progress in any of the tasks. Moreover, each algorithm exhibits comparable behavior under both CTBR and PRT action spaces, with slightly better final performance under PRT for most methods and tasks.
Figure 7: Training curves of single-agent tasks over five seeds.
Figure 8: Training curves of multi-agent cooperative tasks over five seeds.